# CounselBench: A Large-Scale Expert Evaluation and Adversarial Benchmark of Large Language Models in Mental Health Counseling

### v1

@@ 4 CounselBench-ADV: An Adversarial Benchmark for Surfacing LLM Failures in Counseling
Interestingly, the failure modes revealed through CounselBench-ADV are not randomly distributed across models, but exhibit clear family-level patterns. Models within the same family (e.g. LLaMA, Gemini, and Claude) tend to exibit similar distributions across the six failure modes, while GPT-3.5-Turbo and GPT-4 show distinct patterns. This suggests that failure modes remain relatively consistent within model families, but can shift substantially across major model upgrades (openai2024gpt4technicalreport, ).
5 Related Work
A growing body of work has attempted to construct datasets for evaluating counseling quality. Foundational efforts by Pérez-Rosas et al. perez-rosas-etal-ACL2019-good-counselor introduced one of the first annotated datasets for motivational interviews (MI), distinguishing high- and low-quality sessions based on linguistic and behavioral features. Many of these early datasets were built by extracting and transcribing dialogues from actual counseling sessions between patients and counselors (AnnoMI-ICASSP-2022, ; CUEMPATHY-ISCSLP2022, ; SPARTA-HOPE-data, ; PAIR-Dataset-EMNLP-2022, ). Other datasets focused on detecting markers of mental health disorders such as depression or anxiety MOTIVATE-IJCNN ; gratch-etal-2014-DAIC-WoZ-Dataset ; mentalchat16k , annotating student counseling exchanges ConvCounsel-student-counseling-2024 , or capturing specific therapeutic constructs such as empathy MEDIC-MM2023 ; CUEMPATHY-ISCSLP2022 . To mitigate privacy concerns associated with clinical data, some studies turned to social media ALONE-Dataset or summarized clinical notes MEMO-KDD2022 ; others used LLMs to simulate counseling content ict4awe25 ; BOLT-Chiu-Althoff-2024 . Across these works, scholars recognized specific challenges on dataset construction:
Grounding Evaluation in Clinical Expertise
Most prior work adopts evaluation criteria from psychology literature li-etal-2024-understanding-therapeutic ; chiu2024computationalframeworkbehavioralassessment , but rarely involves clinicians in rubric design. As a result, many rely on untrained crowdworkers or subjective heuristics Young_AI_human_response_2024_chi . In contrast, we develop our evaluation protocol collaboratively with clinical psychologists and scale annotation to 100 licensed or trained professionals, ensuring clinical validity and reproducibility.
Counseling Format: Single-turn vs. Multi-turn.
Many existing datasets rely on multi-turn dialogues from real counseling sessions (wang2024clientcenteredassessmentllmtherapists, ; iftikhar2024therapynlptaskpsychologists, ), which, while realistic, introduce challenges for evaluation. For instance, nonverbal cues are lost in transcripts AnnoMI-ICASSP-2022 ; ConvCounsel-student-counseling-2024 ; SPARTA-HOPE-data , and entangling memory, turn-taking, and contextual reasoning can obscure the source of model failures. Moreover, session-specific nuances often hinder generalization. We instead focus on the under-explored single-turn setting, prevalent on public counseling platforms (e.g., CounselChat), where each user question receives a standalone answer. This design facilitates fine-grained, interpretable evaluations and allows direct comparison between models and humans.
LLM-Assisted Annotation and Simulation.
Recent work has leveraged LLMs both to simulate therapy content BOLT-Chiu-Althoff-2024 ; ict4awe25 and to construct artificial patients Scaffolding_Empathy_patient_sim_2025 or therapists Wang_sim_therapists_2024 . LLMs have also been applied as automated evaluators of counseling responses li-etal-2024-understanding-therapeutic ; yosef-etal-2024-assessing ; zhao2024escevalevaluatingemotionsupport , with some studies showing that modern models encode useful therapeutic knowledge ConceptPsy_2025 . However, most of these efforts relied on very limited expert validation. Our work addresses this gap by systematically comparing LLM-generated evaluations against scores from 100 mental health professionals across 2,000 responses. This constitutes the most comprehensive comparison to date between LLM and human judgments in mental health counseling.
6 Limitations, Conclusion, & Future Work
This work introduces a six-dimensional evaluation paradigm for single-turn counseling responses and publishes a corpus of 2,000 expert-annotated evaluations, coupling numeric ratings with in-depth written rationales (CounselBench-EVAL), and 120 clinician-designed adversarial prompts (CounselBench-ADV). To our knowledge, CounselBench is the largest expert-annotated benchmark in this domain. It is the first to encompass both the expertise and the scale of annotators. Our results show that state-of-the-art LLMs outperform community up-voted responses written by online therapists on all quality dimensions, yet our analysis also uncovers persistent pitfalls when LLMs were used either to generate or evaluate counseling responses.
We acknowledge that the current work entails several limitations despite our best efforts. First, although the human-written responses were top-rated on online counseling platforms, they were produced with limited patient information and no in-person contact, which might not represent the best quality of clinician-provided counseling. In fact, clinical psychology literature has long recognized that even well-crafted counseling responses may carry structural biases Smith_counseling_bias_1980 ; clinical-law-review-biases-2002 ; Morrow_Deidan_counseling_bias_1992 ; this highlights the need to diversify comparative standards for evaluating LLM-generated counseling quality. Relatedly, although we aimed to align our expert annotator demographics with real-world trends in the counseling profession, the field itself has long exhibited demographic imbalances.
Alongside addressing these limitations, overcoming current model shortages will require aligning evaluation models more closely with expert standards, for example, through fine-tuning or reinforcement learning with human feedback (RLHF). Expanding the taxonomy of models’ failure modes and integrating them into training and assessment is another pressing direction. Future work could also extend the evaluation paradigm to multi-turn dialogues to see how risks evolve as context accumulates, and replicate the study across languages, clinical domains, and geolocations to test the generality of our current case studies while examining the utility of CounselBench for developing more robust models for clinical use. Together, these efforts will move the field toward LLMs that are not only capable but consistently safe and trustworthy in mental health applications.
@@ Appendix A Survey Design
We utilized Qualtrics to conduct our surveys. We built a template as shown in Figure 5 and then populated surveys for all question and response pairs. Each annotator evaluated five randomly sampled questions, each accompanied by four answers — one from GPT-4, Llama-3.3, Gemini-1.5-Pro, and an online human therapist. For each post-response pair, annotators evaluated across our six metrics, providing rationals and span-level annotations.
Figure 5: Survey interface: annotators read a user post and one response (left) and rate the response on six criteria with Likert scales and text-evidence boxes (right).
@@ Appendix B Annotator Demographic Overview
This section summarizes annotator demographics. Table LABEL:tab:license_degree_summary reports the distribution of counseling licenses and degree types, while Table 5 lists the specialization areas of the 100 mental-health professionals. Figure 6 shows the ethnicity and gender distribution of annotators.
@@
Appendix C Experiment Details
C.1 Response Generation
@@
We use the following prompt to the LLM to generate a response for each user question, and then provide the original user query directly to generate a response.

### v2

@@ 5 CounselBench-ADV: An Adversarial Benchmark for Surfacing LLM Failures
Interestingly, the failure modes revealed by CounselBench-ADV are not randomly distributed across models, but exhibit clear family-level patterns. Models within the same family (e.g., LLaMA, Gemini, and Claude) have similar distributions across the failure modes, while GPT family show distinct patterns. This suggests that failure modes remain relatively consistent within model families, but can shift substantially across major model upgrades (OpenAI et al., 2024; Chen et al., 2023a).
6 Discussion
This work presents CounselBench, a benchmark that evaluates LLMs on open-ended mental health QA. Unlike prior resources that rely on multiple-choice tasks or small-scale expert review, it brings together two complementary components: a large-scale corpus of 2,000 expert evaluations with detailed written rationales (CounselBench-Eval), and a set of 120 adversarial questions written by clinicians to expose failure modes (CounselBench-Adv). Together, they provide both breadth and depth: broad coverage through systematic expert scoring, and targeted probing through adversarial design. Our analysis shows that LLMs can achieve high ratings on several dimensions of quality, yet continue to exhibit recurring weaknesses, including unconstructive feedback, overgeneralization, limited personalization, and unauthorized medical advice. These findings highlight both the progress and the open challenges in deploying LLMs responsibly in mental health QA.
Both components of CounselBench are released as public resources to support future research. For example, CounselBench-Eval can be used to (1) train and validate alignment methods such as reinforcement learning from human feedback, (2) develop automated quality or safety detectors grounded in span-level annotations, and (3) study expert rationales to build critique models or interpretability tools. CounselBench-Adv enables complementary research, such as (1) stress-testing new models to compare robustness under clinically relevant failure modes, (2) incorporating adversarial cases into red-teaming and auditing pipelines, and (3) using these targeted examples as hard negatives for fine-tuning or as feedback signals in multi-agent critique systems. By releasing these resources, we aim to provide a foundation for both rigorous benchmarking and practical model improvement in mental health QA and related high-stakes applications.
@@ Acknowledgments
Ethics Statement
This study was reviewed and approved as exempt research by the Institutional Review Board of University of Southern California (Study ID: UP-25-00022). Questions and clinician responses are sampled from MIT-licensed CounselChat dataset (Bertagnolli, 2020). All identifiers (e.g., names, homepage links) were removed prior to use. Human annotators provided informed consent, including consent for release of de-identified data, and no personally identifiable information was collected.
Reproducibility statement
We release all data and source code. Full hyperparameters, implementation details, random seeds, and prompt templates are in Appendix D, and computational resources are detailed in Appendix LABEL:compute_resouces, to ensure replicability.
@@ Appendix A Survey Design
We utilized Qualtrics to conduct our surveys. We built a template as shown in Figure 4 and then populated surveys for all question and response pairs. Each annotator evaluated five randomly sampled questions, each accompanied by four answers — one from GPT-4, Llama-3.3, Gemini-1.5-Pro, and an human therapist. For each post-response pair, annotators evaluated across our metrics, providing rationals and span-level annotations. At the end of the survey, we collected demographic and professional background data, including gender, ethnicity, freelance capacity, counseling experience, credentials, and areas of specialization.
Figure 4: Survey interface: annotators read a user post and one response (left) and rate the response on criteria with Likert scales and text‐evidence boxes (right).
@@ Appendix B Annotator Demographic Overview
We asked all applicants to provide when they apply: 1) List at least one relevant counseling certifications, licenses, or degrees. 2) Provide a brief overview of their counseling experience.
We further verified the credentials of our 100 expert annotators through multiple (disjoint) methods: (1) license numbers if provided (34 cases), (2) professional profile on platforms such as Psychology Today (38 cases), (3) working and education history on Linkedin (21 cases), (4) listings in workplace or school directories (3 cases), (5) employment and education history on public Upwork profiles (4 cases). We ensured that all annotators had completed counseling-related training or possessed relevant counseling experience.
Table LABEL:tab:license_degree_summary reports the distribution of counseling licenses and degree types, while Table 4 lists the specialization areas of the 100 mental-health professionals. Figure 5 shows the ethnicity and gender distribution of annotators. All data are from background questions in the survey.
@@
Appendix C Annotation Analysis
C.1 Annotation Time and Rationale Length
For quality evaluation, we recorded both the length of the written rationale and time spent per annotator. The median written rationale length was 576.5 words across 20 QA pairs (IQR: 339.75 - 866.5), and the median time spent was 1 hr 22 min (IQR: 52 min - 3 hrs 9 min), indicating substantial engagement.
To examine how time relates to scoring, we binned annotators into quartiles by total time spent (0 – 25%, 25 – 50%, 50 – 75%, 75 – 100%) and compared outcome scores across these groups. To ensure a fair comparison, we focused on the 460 Q&A pairs rated by annotators from all four quartiles. The Table 6 shows the mean score per dimension across quartiles, along with the Kruskal-Wallis $p$-values (for ordinal ratings) and chi-squared test (for categorical variables, i.e. Medical Advice and Factual Consistency because of the option “I am not sure”).
Table 6: Mean metric scores by annotator time quartile (0–25%, 25–50%, 50–75%, 75–100%) for the 460 Q&A pairs rated by annotators from all quartiles. The bottom row reports $p$-values for differences across quartiles (Kruskal–Wallis test for ordinal metrics; chi-squared test for categorical metrics).
Time Quartile Overall Empathy Specificity Medical Advice Factual Consistency Toxicity 0-25% (lowest) 3.29 2.93 3.58 0.13 3.48 1.82 25-50% 3.32 3.32 3.63 0.18 3.29 1.83 50-75% 3.36 3.11 3.77 0.07 3.38 1.83 75-100% (highest) 3.29 2.95 3.64 0.2 3.5 1.5 $p$-value 0.98 0.12 0.61 << .001 << .001 0.08
We found statistically significant associations for Medical Advice and Factual Consistency, which involve complex factual judgments. For example, annotators in the top quartile tended to give more conservative Medical Advice scores (Medical Advice=0.20) than the lowest quartile (Medical Advice=0.13). However, we emphasize that these are associational patterns, not causal effects. Longer time may indicate greater care or, alternatively, more difficult cases for the annotator.
Finally, we assessed whether time correlates with rationale length. The Table 7 shows the average word count per annotator across quartiles. As expected, annotators who spent more time tended to write longer rationales, suggesting greater annotation effort.
Table 7: Average total rationale word count per survey by annotator time quartile (0–25%, 25–50%, 50–75%, 75–100%). Differences across quartiles are significant (by the Kruskal–Wallis test); annotators who spent more time wrote longer rationales on average.
Time Quartile 0-25% (lowest) 25-50% 50-75% 75-100% (highest)
Average Word Count 416.16 604.88 966.36 769.08
$p$-value << .001
C.2 Annotator Counseling Length of Experience
We examined whether counseling experience is associated with rating behavior. We omitted categories with too few annotators, and chose evaluations by annotators from four experience bands (1–3, 3–5, 5–10, and 10+ years). To ensure comparability, we restricted analysis to the 64 question & response pairs that received at least one evaluation from annotators in each band. For each metric, independence between score category and experience band was tested using the chi-squared test. For Medical Advice and Factual Consistency, “I am not sure” responses were excluded; Medical Advice was treated as a binary per-response indicator.
[…]
Table 8: Mean $\pm$ s.d. scores by annotator counseling experience (1–3, 3–5, 5–10, 10+ years) on the 64 question& response pairs rated by all groups. The bottom row reports $p$-values from P tests of independence between score category and experience band; “I am not sure” responses are excluded for Medical Advice and Factual Consistency; Medical Advice is binary (per response).
Overall Empathy Specificity Medical Advice Factual Consistency Toxicity 1-3 years 3.40 $\pm$ 1.24 3.58 $\pm$ 1.14 3.96 $\pm$ 1.13 0.19 $\pm$ 0.39 3.46 $\pm$ 0.74 1.91 $\pm$ 1.20 3-5 years 3.38 $\pm$ 1.24 3.07 $\pm$ 1.41 3.79 $\pm$ 1.26 0.08 $\pm$ 0.27 3.38 $\pm$ 0.68 1.75 $\pm$ 1.20 5-10 years 3.42 $\pm$ 1.30 3.38 $\pm$ 1.30 3.83 $\pm$ 1.20 0.02 $\pm$ 0.15 3.51 $\pm$ 0.91 1.68 $\pm$ 1.24 10+ years 3.30 $\pm$ 1.22 2.97 $\pm$ 1.35 3.92 $\pm$ 1.17 0.08 $\pm$ 0.27 3.37 $\pm$ 0.87 1.96 $\pm$ 1.34 $p$-value 0.90 0.02 0.83 0.04 0.006 0.27
Appendix D Experiment Details
D.1 Response Generation
@@
We use LABEL:prompt:system-prompt to the LLM to generate a response for each user question, and then provide the original user query directly to generate a response.